# TropNNC: Structured Neural Network Compression Using Tropical Geometry

### v1

@@ 7 Experiments
Baselines. We compare our uniform pruning framework with state-of-the-art methods that conduct structured pruning without requiring re-training (or fine-tuning). Specifically, we compare it with the algorithms proposed by Misiakos et al. (2022), ThiNet (Luo et al., 2017), CUP (Duggal et al., 2021) and the simple baselines random and L1 structured pruning. The algorithms presented in (Misiakos et al., 2022) are originally designed for linear layers only. To enable a fair comparison in compressing convolutional layers, we extended their approach using our proposed technique, making them applicable to convolutional layers as well. ThiNet employs a greedy criterion to remove neurons and channels that contribute least to the subsequent layer’s input. For the non-uniform pruning variant of our framework, we compare it with CUP. CUP assigns to each neuron/channel a feature, performs hierarchical clustering based on these features, and replaces each neuron in each cluster by a cluster representative according to a maximum norm criterion. To ensure a fair comparison, we use identical cluster vectors for both frameworks, as defined in CUP, to emphasize the advantages of our tropical geometry-based approach for selecting cluster representatives. Furthermore, the trivial baseline of random structured pruning discards neurons or channels based on a uniform probability distribution, while L1 structured pruning targets those with the smallest L1 norm of their weights or kernels.
Datasets and networks. We evaluate our framework on the MNIST and CIFAR datasets, testing it across various models including simple multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), LeNet (LeCun et al., 1998), AlexNet (Krizhevsky et al., 2012), and VGG (Simonyan and Zisserman, 2015). The non-uniform variant of our algorithm is tested on the CIFAR and ImageNet datasets, with its performance evaluated across models such as VGG, ResNet56-bn, and ResNet18 (He et al., 2016).
@@ 7.2 MNIST and Fashion-MNIST Datasets
To evaluate the performance of our algorithm in compressing linear layers of deeper networks, we applied TropNNC to ”deepNN,” a fully connected neural network with layer sizes $28\times 28,512,256,128$, and $10$. The performance plots are provided in Figures 7(a) and 7(b). As illustrated, TropNNC outperforms both Neural Path K-means from (Misiakos et al., 2022) and ThiNet from (Luo et al., 2017).
To assess the performance of our algorithm in compressing convolutional layers, we applied TropNNC to ”deepCNN2D,” a LeNet-type convolutional neural network with ReLU activations. The performance plots are provided in Figures 7(c) and 7(d). The results demonstrate that TropNNC outperforms Neural Path K-means from (Misiakos et al., 2022) and matches, or even surpasses, ThiNet from (Luo et al., 2017).
@@ 7.3 CIFAR Datasets
In this experiment, we compress AlexNet and VGG trained on the CIFAR-10 and CIFAR-100 datasets to assess the performance of each compression method. Figures 8(a) and 8(b) illustrate the compression of the linear layers of AlexNet on CIFAR-10 and CIFAR-100, respectively. Additionally, Figures 8(c) and 8(d) show the compression of VGG’s convolutional layers for these datasets.
The results indicate that for larger datasets, TropNNC consistently outperforms Neural Path K-means from (Misiakos et al., 2022). Furthermore, it at least matches or even surpasses the performance of ThiNet from (Luo et al., 2017) for the compression of linear layers. These findings highlight the effectiveness of TropNNC in handling more complex and larger-scale data scenarios. For the compression of convolutional layers of VGG, TropNNC matches ThiNet. For the compression of convolutional layers of VGG, we do not present results for iterative TropNNC due to the excessively long execution time required.
@@
Figure 8: Compression of linear and convolutional layers of AlexNet on CIFAR datasets.
@@ 7.4 Non-ReLU Activation
In this experiment, we evaluate the performance of our algorithm in compressing networks that do not use ReLU activations, such as LeNet (LeCun et al., 1998). We conduct experiments on the MNIST and FASHION-MNIST datasets, focusing on compressing the linear and convolutional layers of LeNet. The results are presented in Figures 9(a), 9(b), 9(c) and 9(d). Interestingly, our algorithm demonstrate robust performance even on networks with non-ReLU activations.
@@ 7.5 Non-uniform Pruning
In this section, we evaluate the effectiveness of the non-uniform pruning variant of TropNNC. Our experiments here compress various models, such as VGG on CIFAR-10, ResNet56-bn on CIFAR-10, and ResNet18 on ImageNet. The results are summarized in Tables 4, 5, and 6. We should emphasize that we did not apply any form of fine-tuning or re-training.
Our findings indicate that the method we use for selecting cluster representatives demonstrates a clear advantage. Notably, our approach shows a significant performance improvement for the VGG model, while the benefits are comparatively modest for ResNet models. The substantial advantage observed with VGG remains somewhat unexplained, and we have yet to determine why this effect is less pronounced for ResNet architectures. Further investigation may be required to fully understand these discrepancies and to optimize our approach across different model types.
@@
CUP 0.15 6.24M 0.46G 92.34
TropNNC 0.17 5.90M 0.45G 93.27
CUP 0.20 4.62M 0.42G 67.82
TropNNC 0.22 4.58M 0.41G 92.91
CUP 0.25 3.61M 0.39G 18.10
TropNNC 0.30 3.42M 0.38G 79.84
Table 4: Comparison of CUP and TropNNC accuracy across different pruning thresholds on CIFAR10, VGG.
@@
CUP 0.50 0.83M 0.22G 87.10
TropNNC-3it. 0.50 0.83M 0.22G 88.34
CUP 0.55 0.82M 0.21G 82.18
TropNNC-3it. 0.55 0.82M 0.21G 83.32
CUP 0.60 0.79M 0.20G 59.97
TropNNC-3it. 0.60 0.79M 0.20G 64.16
Table 5: Comparison of CUP and TropNNC (3 iterations) accuracy across different pruning thresholds on CIFAR10, ResNet56-bn.
@@
Original - 11.68M 3.63G 69.10
CUP 0.60 11.65M 3.55G 62.30
TropNNC 0.60 11.65M 3.55G 64.80
CUP 0.65 11.60M 3.48G 54.10
TropNNC 0.65 11.60M 3.48G 56.60
CUP 0.70 11.43M 3.32G 29.20
TropNNC 0.70 11.43M 3.32G 33.20
@@ 8 Conclusions and Future work
We improved upon the theoretical results of Misiakos et al. (2022) by using the (usual) Hausdorff distance instead of its discrete counterpart, to obtain a tighter bound for the approximation of tropical polynomials. Leveraging this enhancement we refined the choice of compressed node’s weights that correspond to the K-means clusters from (Misiakos et al., 2022). Our novel algorithm, TropNNC, applied to linear and convolutional layers of deep neural networks outperforms Neural Path K-means Misiakos et al. (2022), and manages to match or even surpass the performance of ThiNet, particularly for the compression of linear layers. In non-uniform pruning it performs better than CUP, with significant improvement in the case of the VGG network architecture. These achievements underscore the potential of tropical geometry in the realm of neural network compression. One promising direction for further investigation could involve optimizing the approximation techniques by considering the upper hulls of the network’s tropical polynomials.
@@ A.1 Proof of Lemma 2
$=\sum_{i=1}^{k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq\max_{i=1,\ldots,k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq p(\mathbf{x})$
@@ A.2 Proof of Lemma 8
$f(\mathbf{u})=\operatorname{dist}(\mathbf{u},\tilde{P})=\min_{\tilde{\mathbf{u}}\in\tilde{P}}\|\mathbf{u}-\tilde{\mathbf{u}}\|$
@@ A.3 Proof of Proposition 10 and Theorem 11
$\leq\sum_{j=1}^{m}\left(\lvert p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x})\rvert+\lvert q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x})\rvert\right)$
@@
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right),$
@@ A.4 Proof of Proposition 12
$v(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x}),\quad\tilde{v}(\mathbf{x})=\tilde{p}(\mathbf{x})-\tilde{q}(\mathbf{x})$
@@
$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq H(P,\tilde{P})+H(Q,\tilde{Q})$
@@
Let $I_{+}\subset[n]$ and $I_{-}\subset[n]$ be the sets of positive and negative generators, respectively. First, we deal with $H(P,\tilde{P})$. Notice that $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is the sum of generators of some clusters of $P$. Thus $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is a vertex of $P$. Hence,
@@
Let $I_{k}\subset[n]$ be the set of generators that belong to cluster $k$. Let $C_{+}\subset[K]$ be the set of clusters of positive generators, and $C_{-}\subset[K]$ the set of clusters of negative generators.
Consider any vertex $\mathbf{u}$ of $P$. This vertex can be written as the sum of generators $c_{i}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{+}^{\prime}\subset I_{+}$. Thus,
$\mathbf{u}=\sum_{i\in I_{+}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i})$
For every generator $c_{i}(\mathbf{a}_{i}^{T},b_{i})$ that belongs to cluster $k$, define
$x_{i}=\arg\min_{x}\left\{c_{i}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\}$
@@
Define $I_{k}^{\prime}=I_{k}\cap I_{+}^{\prime}$. For every cluster $k$, define
$\tilde{x}_{k}=\sum_{i\in I_{k}^{\prime}}x_{i}$
@@
Thus, for every cluster $k$, the point $\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k}^{T})$ lies inside the segment $[\mathbf{0},\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k}^{T})]$ and thus belongs to $\tilde{P}$.
For the vertex $\mathbf{u}$, we choose to compare it with the vertex
$\sum_{k\in C_{+}}\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k}^{T})\in\tilde{P}.$
@@
$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x_{i}|I_{k}|(c_{i}(\mathbf{a}_{i}^{T},b_{i})+\epsilon_{i})\right\|$
@@
$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\left\{|1-x|I_{k}||\cdot\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|+|x|I_{k}||\cdot\left\|\epsilon_{i}\right\|\right\}$
@@
$=\sum_{i\in I_{+}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
@@
$\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P})\leq\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
@@
$=\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
@@
$H(Q,\tilde{Q})\leq\sum_{i\in I_{-}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$
@@
$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq\sum_{i\in I}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}$

### v2

@@ 7 Experiments
Baselines. We compare our framework with state-of-the-art methods that conduct structured pruning without requiring re-training (or fine-tuning). Specifically, we compare it with the algorithms proposed by Misiakos et al. (2022), ThiNet (Luo et al., 2017), CUP (Duggal et al., 2021) and the simple baselines random and L1 structured pruning. The algorithms presented in (Misiakos et al., 2022) are originally designed for linear layers only. To enable a fair comparison in compressing convolutional layers, we extended their approach using our proposed technique, making them applicable to convolutional layers as well. ThiNet employs a greedy criterion to remove neurons and channels that contribute least to the subsequent layer’s input. For the non-uniform pruning variant of our framework, we compare it with CUP. CUP assigns to each neuron/channel a feature, performs hierarchical clustering based on these features, and replaces each neuron/channel in each cluster by a cluster representative according to a maximum norm criterion. As explained, our proposed algorithm enhances all three steps of CUP. Through an ablation study, we found that each improvement is essential for achieving a competitive advantage over CUP, indicating that these enhancements complement each other. In the presented experiments, we compare CUP exclusively with the fully enhanced version of TropNNC, incorporating all improvements. Furthermore, the trivial baseline of random structured pruning discards neurons or channels based on a uniform probability distribution, while L1 structured pruning targets those with the smallest L1 norm of their weights or kernels.
Datasets and networks. We evaluate our framework on the MNIST and CIFAR datasets, testing it across various models including simple multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), LeNet (LeCun et al., 1998), AlexNet (Krizhevsky et al., 2012), and VGG (Simonyan and Zisserman, 2015). The non-uniform variant of our algorithm is tested on the CIFAR and ImageNet datasets, with its performance evaluated across models such as VGG, ResNet56, and ResNet18 (He et al., 2016).
@@ 7.2 MNIST and Fashion-MNIST Datasets
To evaluate the performance of our algorithm in compressing linear layers of deeper networks, we applied TropNNC to ”deepNN”, a fully connected neural network with layer sizes $28\times 28,512,256,128$, and $10$. The performance plots are provided in Figures 7(a) and 7(b). As illustrated, TropNNC outperforms both Neural Path K-means from (Misiakos et al., 2022) and ThiNet from (Luo et al., 2017).
To assess the performance of our algorithm in compressing convolutional layers, we applied TropNNC to ”deepCNN2D”, a LeNet-type convolutional neural network with ReLU activations. The performance plots are provided in Figures 7(c) and 7(d). The results demonstrate that TropNNC outperforms Neural Path K-means from (Misiakos et al., 2022) and matches, or even surpasses, ThiNet from (Luo et al., 2017).
@@ 7.3 CIFAR Datasets
In this experiment, we compress AlexNet and VGG trained on the CIFAR-10 and CIFAR-100 datasets to assess the performance of each compression method. Figures 8(a) and 8(b) illustrate the compression of the linear layers of AlexNet on CIFAR-10 and CIFAR-100, respectively. Additionally, Figures 8(c) and 8(d) show the compression of VGG’s convolutional layers for these datasets. For VGG, we applied compression to the fused networks without batch normalization, as we found that ThiNet’s performance on the original networks was practically identical to its performance on the fused networks.
The results indicate that for larger datasets, TropNNC consistently outperforms Neural Path K-means from (Misiakos et al., 2022). Furthermore, it at least matches or even surpasses the performance of ThiNet from (Luo et al., 2017) for the compression of linear layers. These findings highlight the effectiveness of TropNNC in handling more complex and larger-scale data scenarios. For the compression of convolutional layers of VGG, TropNNC matches ThiNet. For the compression of convolutional layers of VGG, we do not present results for iterative TropNNC because it had almost identical performance to the non-iterative algorithm.
@@
Figure 8: Compression of linear layers of AlexNet and convolutional layers of VGG on CIFAR datasets.
@@ 7.4 Non-ReLU Activation
In this experiment, we evaluate the performance of our algorithm in compressing networks that do not use ReLU activations, such as LeNet (LeCun et al., 1998). We conduct experiments on the MNIST and FASHION-MNIST datasets, focusing on compressing the linear and convolutional layers of LeNet. The results are presented in Figures 9(a), 9(b), 9(c) and 9(d). Interestingly, our algorithm demonstrates robust performance even on networks with non-ReLU activations.
@@ 7.5 Non-uniform Pruning
In this section, we evaluate the effectiveness of the non-uniform variant of TropNNC. Our experiments here compress various models, such as VGG on CIFAR-10, ResNet56 on CIFAR-10, and ResNet18 on ImageNet. The results are summarized in Tables 4, 5, 6, and 7. We should emphasize that we did not apply any form of fine-tuning or re-training. In Table 4 we present the compression of fused models, i.e. models whose batch normalization has been fused into its preceding convolutional or linear layer, as explained in a previous section. We noticed that the performance of CUP was better on the original networks with batch normalization. In Tables 5, 6 and 7 we compare our performance with the performance of CUP compressing the original networks. For our method only, we applied batch normalization fusion following the second approach described earlier—fusing the batch normalization layer with the preceding layer on a layer-by-layer basis, followed by compression of the fused layer before moving on with the next layer. We observed that our method performed significantly better when the clustering vectors were generated using the pre-fusion weights/kernels of the layers. While the reason for this improvement remains unclear, it consistently led to better performance in our experiments.
Our findings indicate that our method demonstrates a clear advantage. Notably, our approach shows a significant performance improvement for the VGG model, while the benefits are comparatively modest for ResNet models. The substantial advantage observed with VGG remains somewhat unexplained, and we have yet to determine why this effect is less pronounced for ResNet architectures. Further investigation may be required to fully understand these discrepancies and to optimize our approach across different model types.
We also found that the first variant of non-uniform TropNNC excelled at reducing the overall network size but was less effective at minimizing inference operations. In contrast, the second variant performed well in both tasks, outperforming CUP. Upon further analysis, we concluded that the first variant tends to focus more aggressively on the final layers, where parameter count is high due to the increased number of channels, but the number of operations is lower because of smaller image sizes. Meanwhile, the second variant, like CUP, also targets the initial layers, where fewer parameters are present, but a greater number of operations is required.
Model Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
VGG Original - 14.72M 0.63G 93.64
CIFAR10 CUP 0.15 4.70M 0.40G 56.22
TropNNC (v2) 1.12 4.70M 0.33G 90.82
ResNet56 Original - 0.85M 0.25G 93.67
CIFAR10 CUP 0.55 0.79M 0.18G 45.89
TropNNC (v2) 1.2 0.65M 0.18G 76.50
ResNet18 Original - 11.68M 3.63G 69.10
ImageNet CUP 0.6 11.62M 3.50G 58.20
TropNNC (v2) 1 11.60M 3.47G 62.60
Table 4: Comparison of CUP and TropNNC (variant 2) accuracy across different pruning thresholds for fused models without batch normalization.
@@
CUP 0.15 6.24M 0.46G 92.41
TropNNC (v1) 0.006 5.70M 0.44G 93.58
CUP 0.20 4.62M 0.42G 67.91
TropNNC (v1) 0.0085 4.32M 0.41G 93.46
CUP 0.25 3.61M 0.39G 16.02
TropNNC (v1) 0.011 3.53M 0.39G 93.21
Table 5: Comparison of CUP and TropNNC (variant 1) accuracy across different pruning thresholds on CIFAR10, VGG.
@@
CUP 0.55 0.72M 0.20G 87.32
TropNNC (v2) 1.17 0.71M 0.19G 90.53
CUP 0.6 0.63M 0.17G 78.83
TropNNC (v2) 1.23 0.63M 0.17G 87.04
CUP 0.65 0.54M 0.15G 71.48
TropNNC (v2) 1.3 0.54M 0.15G 73.03
TropNNC (v1) 0.067 0.49M 0.19G 82.70
Table 6: Comparison of CUP and TropNNC accuracy across different pruning thresholds on CIFAR10, ResNet56.
@@
Original - 11.69M 3.64G 70.10
CUP 0.5 11.66M 3.58G 66.60
TropNNC (v2) 1.1 11.64M 3.44G 68.40
CUP 0.6 11.49M 3.38G 53.20
TropNNC (v2) 1.2 11.27M 3.07G 57.90
CUP 0.65 11.17M 3.15G 29.20
TropNNC (v2) 1.25 10.71M 2.75G 41.20
TropNNC (v1) 0.021 9.87M 3.46G 58.10
@@ 8 Conclusions and Future work
We improved upon the theoretical results of Misiakos et al. (2022) by using the (usual) Hausdorff distance instead of its discrete counterpart, to obtain a tighter bound for the approximation of tropical polynomials. Leveraging this enhancement we refined the choice of compressed node’s weights that correspond to the K-means clusters from (Misiakos et al., 2022). Our novel algorithm, TropNNC, applied to linear and convolutional layers of deep neural networks outperforms Neural Path K-means of Misiakos et al. (2022), and manages to match or even surpass the performance of ThiNet, particularly for the compression of linear layers. In non-uniform pruning it performs better than CUP, with significant improvement in the case of the VGG network architecture. These achievements underscore the potential of tropical geometry in the realm of neural network compression. One promising direction for further investigation could involve optimizing the approximation techniques by considering the upper hulls of the network’s tropical polynomials.
@@ A.1 Proof of Lemma 2
$=\sum_{i=1}^{k}\lambda_{i}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq\max_{i=1,\ldots,k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq p(\mathbf{x}).$
@@ A.2 Proof of Lemma 8
$f(\mathbf{u})=\operatorname{dist}(\mathbf{u},\tilde{P})=\inf_{\tilde{\mathbf{u}}\in\tilde{P}}\|\mathbf{u}-\tilde{\mathbf{u}}\|$
@@ A.3 Proof of Proposition 10 and Theorem 11
$\leq\sum_{j=1}^{m}\left(\lvert p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x})\rvert+\lvert q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x})\rvert\right).$
@@
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right).$
@@ A.4 Proof of Proposition 12
$v(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x}),\quad\tilde{v}(\mathbf{x})=\tilde{p}(\mathbf{x})-\tilde{q}(\mathbf{x}).$
@@
$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq H(P,\tilde{P})+H(Q,\tilde{Q}).$
@@
Let $I_{+}\subseteq[n]$ and $I_{-}\subseteq[n]$ be the sets of positive and negative generators, respectively. First, we deal with $H(P,\tilde{P})$. Notice that $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is the sum of generators of some clusters of $P$. Thus $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is a vertex of $P$. Hence,
@@
Let $I_{k}\subseteq[n]$ be the set of generators that belong to cluster $k$. Let $C_{+}\subseteq[K]$ be the set of clusters of positive generators, and $C_{-}\subseteq[K]$ the set of clusters of negative generators.
Consider any vertex $\mathbf{u}$ of $P$. This vertex can be written as the sum of generators $c_{i}(\mathbf{a}_{i}^{T},b_{i})$, for some subset $I_{+}^{\prime}\subseteq I_{+}$. Thus,
$\mathbf{u}=\sum_{i\in I_{+}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i}).$
For every positive generator $c_{i}(\mathbf{a}_{i}^{T},b_{i})$ that belongs to cluster $k$, define
$x_{i}=\arg\min_{x}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|,$
@@
For every cluster $k\in C_{+}$ define $I_{k}^{\prime}=I_{k}\cap I_{+}^{\prime}$ and
$\tilde{x}_{k}=\sum_{i\in I_{k}^{\prime}}x_{i}.$
@@
Thus, for every cluster $k\in C_{+}$, the point $\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\mathbf{0},\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})]$ and thus belongs to $\tilde{P}$.
For the vertex $\mathbf{u}$, we choose to compare it with the point
$\sum_{k\in C_{+}}\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\in\tilde{P}.$
@@
$=\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})-x\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
@@
$\leq\sum_{k\in C_{+}}\sum_{i\in I_{k}^{\prime}}\min_{x}\left\{|1-x|I_{k}||\cdot\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|+|x|I_{k}||\cdot\left\|\epsilon_{i}\right\|\right\}$
@@
$=\sum_{i\in I_{+}^{\prime}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},$
@@
$\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P})\leq\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
@@
$=\sum_{i\in I_{+}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
@@
$H(Q,\tilde{Q})\leq\sum_{i\in I_{-}}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\}.$
@@
$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq\sum_{i\in I}\min\left\{\left\|c_{i}(\mathbf{a}_{i}^{T},b_{i})\right\|,\delta_{\max}\right\},$